HUBBLE SPACE TELESCOPE DAILY REPORT #931 PERIOD COVERED: 0800 EST Daily Status Report as of 198/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 5056 (High Speed Astrometry - A Search for Planetary Companions to Low Mass) The Fine Guidance Sensors were used to make astrometric observations of Bernard's star to look for the presence of a Jupiter sized planet in orbit around the target star from the position perturbations caused by the planet. The astrometric observations were monitored in the OSS, and no problems were seen in the telemetry. 1.2 Completed Five Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to complete one ultraviolet snapshot observations fo the nearby galaxy UGC-12113. The exposure lasted for ten minutes and used filter F220W. The observation was reported to be satisfactory. 1.3 Completed Five Sets of FOS/RD 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3 : Cycle 3 High) The Faint Object Spectrograph was used to make 2 observation of quasar 1306+3021 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. Both observations were received and displayed. The spectrum from the following observations were reported to look fine. 1.4 Completed WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make two Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. Both observations were received. One observation was reported to be saturated, and the other was okay. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed Three Sets of FOC/96 4608 (Multi Timescale Multispectral Observation of the Jovian Aurora Cycle 3) The Faint Object Camera was used to make observations of the Jovian Aurorae using a range of filters to study the intensity of the aurorae, and their wavelength dependence. A dark calibration image was successfully completed. The OSS verified that the correct filter was in place before each set of observations. All of the observations were reported to look excellent. 1.7 Completed PC/WFC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. daises and measles), and the possible return of quantum efficiency hysteresis. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.8 Completed Two Sets of FOS/BL 4548 (Search for Evidence of Active Ice Volcanism on Europa Cycle 3 High) The Faint Object Spectrograph was used to make three observations of the surface of Jupiter's moon Europa looking for water volcanoes. Some of the ejected water will be photo-dissociated to produce OH, which should be detected by the emission band at 3090 A0. All three observations were reported to look satisfactory. 1.9 Completed FOC/96 4969 (Search for QSOs suitable for Subsequent Observation of He II) The Faint Object Camera was used to make an ultraviolet image of the quasi-stellar object 0234+013 to measure the levels of its ultraviolet flux to see if this object is suitable for subsequent observation with the FOS and GHRS. They are looking for absorption features of Helium II at 304 A0. The observation was executed as scheduled, and the image was reported to be satisfactory. 1.10 Completed FOC/96 4838 (Imaging of Elliptical Galaxies Cycle 3 Med.) The Faint Object Camera (f/96) was used to make two observations of the elliptical galaxy NGC-315 to locate precisely its center, and to measure if its center is bright enough for future spectroscopic observations. Both observations were received and displayed in the OSS. They were reported to look okay. 1.11 Completed WFC 4849 (GTO High Latitude Parallel Program Cycle 3 High) The Wide Field Camera was used to make high latitude sky survey observations taken in parallel with the prime science observations. Two observations were received and displayed. They both were reported to look okay. 1.12 Completed FOS/BL 4396 (QSO Absorption System Snapshot Survey Part 1 of 3: Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make two spectrographic observations of quasar 0151+0433 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The observations were received in real time. The peakup had maximum counts in the first group, and the spectrum was weak. The baseline guide star acquisition failed to get fine lock so the observation was make in course track. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 22 Successful : 22 At 199/2037Z Fine Lock Failed to Course Track during LOS. Subsequent observations were okay. Scheduled Reacquisitions : 12 Successful : 12 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 24 Successful : 24 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.2 Complete FOS/RD 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3 : Cycle 3 High) 3.3 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.4 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.5 Complete FGS 5060 (Linking HIPPARCOS to the Extragalactic Reference Frame Part 2 of 6) 3.6 Complete FOS/BL 4976 (UV Rotation Light Curves of Pluto, and Charon's UV Spectrum - Split) 3.7 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.8 Complete WFC/PC 4109 (WF/PC Cycle Calibration: Rapid Internal Monitor) 3.9 Complete PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) 3.10 Complete FOC/96 4353 (Hubble s Space Telescope Imaging of IO: Cycle 3 High) 3.11 Complete PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) 3.12 Complete PC 3458 (Stellar Population Gradiants in Post-Core- Collapse Globular Clusters) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-